Listed below are the top 10 annual salaries​ (in millions of​ dollars) of TV personalities. Find the​ range, variance, and standard deviation for the sample data. Given that these are the top 10​ salaries, do we know anything about the variation of salaries of TV personalities in​ general? 42    41    39   33   23    20     18    16    15.5     14.7
Answer #1

Solution :

Maximum = 42

Minimum = 14.7

Range = maximum - minimum = 42 - 14.7 = 27.3

n = 10

= = 262.2 / 10 = 26.22

variance = s2 =  (x - )2 / n - 1

= 1145.456 / 9

variance = 127.2729

Standard deviation = 11.2815